Question 11 pts

Oranges b) Oranges tons per year tons per year) 28 Brazil Argentina 16 14 0 Apples (tons per year) Apples (tons per year)

What is the opportunity cost of 1 ton of apples for the nations of Argentina and Brazil, respectively?

4 tons of oranges and 2 tons of oranges
2.5 tons of oranges and .4 tons of oranges
.25 tons of oranges and .5 tons of oranges
2 tons of oranges and 4 tons of oranges
Answer #1

Answer
The opportunity cost of a good= maximum production of other good /maximum production of the good
The opportunity cost of Apple for Argentina=28/14=2 oranges
The opportunity cost of apple for Brazil =16/4=4 oranges
Option fourth

?